A survey was conducted among 40 tourists. The results were 28 visited Antigua 30 visited Barbados and 3x visited both Antigua and Barbados . X visited neither Antigua or Barbados.
1. Write an expression in x To represent the total number of tourist in the survey ?
2. Calculate the value of x

Answer by stanbon(75887)

A survey was conducted among 40 tourists. The results were 28 visited Antigua 30 visited Barbados and 3x visited both Antigua and Barbados . X visited neither Antigua or Barbados.
1. Write an expression in x To represent the total number of tourist in the survey ?
28 + 30 -3x + x = 40
-----------------------------
2. Calculate the value of x
-2x = 40-58
-2x = -18
x = 9
==================
Cheers,
Stan H.
==================